OpenStudy (anonymous):

PLEASE HELP AND EXPLAIN! FOR A MEDAL! 19=h/3-8

OpenStudy (midhun.madhu1987):

Solve for h?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve for H sooo Bring the 8 over 19 +8=27 27=h/3 Now the opposite of multiplying 27x3=H 81+h Now check 19=81/3-8

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Yes, assuming @roz_mar23 meant "81=h", not "81+h". The point is to isolate the term with the h term and then multiply out to get the final answer.

OpenStudy (midhun.madhu1987):